A motel clerk counts his $1 and $10 bills at the end of the day. He finds that he has a total of 60 bills having a combined monetary value of $195. Find the number of bills of each denomination that he has.

Answer by addingup(3677)

He has x 1 dollar bills
He has 60-x 10 dollar bills
The bills total 195:
1x+10(60-x) = 195
1x+600-10x = 195
-9x = -405 Divide both sides by -9, and remember -/- = + so when you divide you can ignore the - because the answer is going to be positive.
x = 45 This is how many 1 dollar bills he has. And:
60-45 = 15 he has 15 10 dollar bills
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
45 * 1 = 45
15 * 10 = 150
45 + 150 = 195 Correct